OpenStudy (shadowlegendx):

If x = ba and y = con What is x + y = ??? What is x(x) + y(y) = ??? EDIT: This isnʻt a serious question. Answer with your imagination :P

OpenStudy (shadowlegendx):

This requires some knowledge of BaCoNoLoGy

OpenStudy (anonymous):

ba+con

OpenStudy (anonymous):

replace x with what it equals

OpenStudy (anonymous):

so if x = ba and y = con ba + con

OpenStudy (anonymous):

bacon

OpenStudy (anonymous):

ba times ba + con times con

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\frac{{\rm sin~}x}{\rm n} =\frac{{\rm si\bcancel{n}~}x}{\rm \bcancel{n}}={\rm six}=6 }\)

OpenStudy (shadowlegendx):

^lel

OpenStudy (defiance):

This is a very nice question.

OpenStudy (defiance):

Therefore, I can not help.

OpenStudy (anonymous):

bacon and eggs?